OpenStudy (anonymous):

What is the solution to the equation 7 –3x ≈ 9 ?

OpenStudy (abb0t):

approximately? Lol

OpenStudy (anonymous):

yea lol these are the choices x = 0.376 x = 0.295 x = –0.295 x = –0.376

OpenStudy (anonymous):

ohh it is just regular problem.

OpenStudy (anonymous):

can you guys help or ?

OpenStudy (anonymous):

sure you have the equation 7 –3x ≈ 9 right? you must first subtract 7 from both sides like this 7 - 7 - 3x = 9 -7 3x = 2 x = 2/3

OpenStudy (anonymous):

oh snap its 7^(-3x) ≈ 9